OpenStudy (anonymous):

A boy walks a certain distance at 9km/hr. He find if he increase his speed by 1/3, he would have save 1 1/3 hours. Find the distance he walk? www.teachon.tk

OpenStudy (perl):

d = 9t d = ( 9 + 1/3*9) ( t - 4/3)

OpenStudy (perl):

I get t = 16/3

OpenStudy (perl):

so d = 9*(16/3) = 48 km

OpenStudy (callisto):

Let t be the original time taken s = x/t, x=st Therefore 9t = 9 [1+(1/3)] (t- 1 1/3) 9t = 12t -16 t = 16/3 Distance travelled = st = 9 (16/3) = 3 x 16 = 48 km Sorry i'm slow

OpenStudy (anonymous):

thanks dear

OpenStudy (callisto):

* the distance travelled is the same so s1t1 = s2t2